It has been known that a fixed dose of a certain drug results to an



average increase of pulse rate by at least 12 beats per minute with



a standard deviation of 5. A group of 20 patients given the same



dose showed the following increases: 15, 12, 10, 8, 14, 15, 16, 11,



7, 13, 9, 10, 12, 11, 9, 10, 17, 14, 15, 7. Is there proof to show that



this group has a lower average increase of pulse rate than the ones



in general? Use 0.05 level of significance

1
Expert's answer
2021-11-19T00:07:15-0500
"z=\frac{11.75-12}{\frac{5}{\sqrt{20}}}=-0.22"

It is greater than negative critical value.  There is not proof to show that

this group has a lower average increase of pulse rate than the ones in general at 0.05 level of significance.